OpenStudy (anonymous):

Find the geometric mean of 48 and 3.

OpenStudy (turingtest):

the geometric mean of two numbers, \(a\) and \(b\) is\[GM=\sqrt{a\times b}\]

OpenStudy (mertsj):

\[\frac{48}{x}=\frac{x}{3}\]

OpenStudy (anonymous):

Let the geometric sequence be : 48, x, 3 (As we are to find mean, so x will lie between 48 and 3 So, how to find common ratio in Geometric progression: r = Second term/First term = Third term/Second term So, x/48 = 3/x \[x^2 = 48 \times 3\] Multiply it and take the square root you will have your answer...